# A Perspective on Foundation Models for the Electric Power Grid

### v1

@@
Abstract
Foundation models (FMs) currently dominate news headlines. They employ advanced deep learning architectures to extract structural information autonomously from vast datasets through self-supervision. The resulting rich representations of complex systems and dynamics can be applied to many downstream applications. Therefore, FMs can find uses in electric power grids, challenged by the energy transition and climate change. In this paper, we call for the development of, and state why we believe in, the potential of FMs for electric grids. We highlight their strengths and weaknesses amidst the challenges of a changing grid. We argue that an FM learning from diverse grid data and topologies could unlock transformative capabilities, pioneering a new approach in leveraging AI to redefine how we manage complexity and uncertainty in the electric grid. Finally, we discuss a power grid FM concept, namely GridFM, based on graph neural networks and show how different downstream tasks benefit.
keywords:
@@ Abstract
footnotetext: § Lead contact: hehamann.hh@gmail.com or hendrikh@us.ibm.comfootnotetext: * These authors contributed equally to this work as major contributors.footnotetext: † These authors contributed equally to this work as significant contributors.footnotetext: ‡ These authors contributed equally to this work.
@@ 1 Introduction
Figure 1: The energy transition, aging infrastructure, cybersecurity challenges, and climate change greatly increase complexity and uncertainty in operating, controlling, and planning the power grid. This is creating a widening gap between existing computational capabilities and the evolving needs of the electric power industry. AC PF denotes alternating current power flow; DC PF denotes direct current power flow.
@@
In this context, it is clear that the operation, control, and planning of power systems will soon be pushed to their limits. As depicted in Figure 1, new computational methods and approaches are needed, capable of better tackling the challenges presented by increased uncertainty and complexity. Machine learning (ML) and artificial intelligence (AI) methods have shown promise for such purposes in a wide spectrum of industries, with significant breakthroughs in computer vision [8], natural language processing [9], and intelligent control [10]. Such approaches rely on statistical algorithms to learn from massive data, often enabling them to execute tasks (e.g., regression, classification, detection, segmentation, forecasting, generation) without human supervision.
@@
Yet, despite significant progress in the application of AI/ML methods to power systems [11, 12, 13], wider adoption has been impeded, in particular, by a lack of readily available training data and the limited transferability of developed models to adjacent applications [14]. This is where, we argue, Foundation models (FM) have an important role to play. FMs are advanced AI models developed through self-supervised learning, most often based on transformer architectures, that generalize across various tasks after initial training on large datasets, enabling efficient adaptation to specific applications with minimal labeled data. Text-based FMs like GPT-4 [15], encoding deep relationships in text, have enabled a wide range of new applications. Similarly, we anticipate that a multi-modal FM for power grids, trained on the diverse data that characterizes the power grid, will enable a broad set of new use cases.
@@
To enable this, we advocate in this article for developing multi-modal FMs for power grids: what we refer to as GridFMs. We explain why we believe in their potential, highlighting their anticipated strengths and weaknesses in light of the challenges posed by the evolving power grid. We develop a concrete picture of what we envision a GridFM to be and how we expect it to pioneer a new approach in leveraging AI to manage complexity and uncertainty in the power grid.
@@
In the following, we first highlight, in Section 2, the looming power grid challenges associated with the energy transition, which are further aggravated by aging infrastructure, swelling cybersecurity, and climate change concerns. We show how, by increasing the complexity and uncertainty of the power grid, these drivers are creating a significant gap in computational capabilities, which we propose to overcome with FMs. We then review in Section 3 the strengths and weaknesses of FMs and propose in Section 4 an implementation of a GridFM that we argue will enable exciting, currently unattainable use cases as downstream tasks.
@@ 2.1 Distributed and renewable energy sources
While a single coal plant can generate hundreds of megawatts predictably and reliably, large numbers of weather-dependent renewable energy resources like wind and solar are required to produce the same power output. Current hourly and daily forecast methods still show up to 3% and up to 10% forecast error, respectively [16]. Additionally, distributed energy resources (DERs) like battery storage and electric vehicles are operated by independent power management systems and are driven by user patterns with intrinsic uncertainties. Dealing with the added uncertainty from reverse DER power flows, multipoint injections, and weather-induced fluctuations requires new voltage regulation methods, protection schemes, and advanced (optimal) power flow analysis. The numbers of controllable and dispatchable devices, as well as balancing nodes are expected to grow exponentially [17] with additional thousands to millions to be added in the near future. Managing the associated variability and uncertainty will require advanced stochastic optimization, scenario reduction, and sampling-based approaches, which are computationally more intensive than currently used deterministic methods, especially for real-time demand-supply balancing [18]. Furthermore, as the net load of DER-rich distribution branches seen by transmission operators becomes more variable and less predictable, effective grid management and planning will become impossible without effective co-simulation frameworks that integrate transmission and distribution analysis. Yet the computational cost of such simulations, at least with existing technologies, is extremely high.
@@ 2.2 Everything inverter
For the future power grid, inverters will play a crucial role in maintaining low cost at the grid’s edge. However, with increased inverter proliferation, established control schemes and grid stability are challenged, while a more deregulated environment is created. With more inverters connected, grid inertia – which stabilizes phase, voltage, and frequency during transient events and which is mostly tied to large synchronous generators – decreases [19], unless more grid-forming converters are introduced. As inverters are driven by software, they can quickly adjust power or disconnect, causing additional issues like sub-synchronous resonances and even trip large generators [20].
Despite the increasing adoption of standards [21], inverters remain hard to predict and model, unlike physics-driven synchronous generators. Quasi-static phasor simulations are used to assess optimal power flow. For grid stability assessment, however, computationally expensive electro-magnetic transient (EMT) simulations [22] are required. They run mostly on traditional central processing units (CPUs), which are flexible for many algorithms but inferior in compute performance to recent graphics processing units (GPUs), which are just starting to get adopted in the industry, as, e.g., in [23]. Also, CPU systems do not always scale well for large EMT simulations [24]. Grid operations, however, would require EMT simulations to run in real-time and for large sections of the grid, including distribution grid sections. To further account for inherent network uncertainties and those introduced by inverter-based resources (IBRs), stochastic ensembles of the problems would need to be simulated. Altogether, such an approach is computationally expensive and, with current approaches, impractical [25]. Thus, the community mostly relies on offline simulators like PS CAD [26] or ParaEMT [27] while also exploring physics-informed neural networks to reduce EMT simulation time [28].
@@ 2.3 Changes to demand and weather patterns
Long-term models like the Regional Energy Deployment System [35] are used for resource, infrastructure, and strategic planning. Uncertainty and variability at longer time scales are significantly driven by climate policy, technology adoption, and regulatory changes. Instances of this include the combustion engine phase-out by 2035 in Europe – which resulted in a massive increase in the long-term electricity demand forecast – or the US Federal Energy Regulatory Commission’s rules on planning and covering the cost of the power grid [36]. To improve model accuracy, additional information that captures demographic changes and changes in the distribution of wealth, which lead to different societal and individual behaviors (e.g., heating, cooling, electromobility), as well as climate change-induced environmental shifts must be integrated. Therefore, the capacity to combine heterogeneous data sources (including information in textual format) is a fundamental requirement to improve predictability.
@@ 2.4 Aging infrastructure
Moreover, the lowering of grid inertia through IBR integration (see Section 2.2) is further amplified by the retirement of old synchronous generators. The more frequent use of synchronous generators for regulating purpose bring them more often in operating zones where mechanical or thermal stresses are significant, and thus accelerate the aging of the generators. IBR related switching transients also disrupt existing protection relay schemes and increase grid behavioral uncertainty.
@@ 3 Foundation models
Having identified the growing gaps between computational capabilities and industry needs amidst the energy transition, we now introduce FMs, which is the focus of this article, laying the groundwork for our vision of GridFM.
@@ 3.1 The foundation model approach
To understand the FM approach, it helps to review the evolution of AI and ML as depicted in Figure 2. Early hand-crafted symbolic expert systems, from the 1960s onwards, were limited and brittle. The 90s brought big data and general-purpose ML that, however, depend on task-specific hand-crafted features. In 2012, deep learning, enabled by increased compute performance, started to disrupt the AI world [45], but remained task-specific and still required large annotated datasets. Recently, FMs have emerged. They learn from data through self-supervision [46] and have proven to generalize across many applications. Creating an FM is a three-step process, as shown in Figure 3:
@@
    Pre-training: Developing a general-purpose representation FM through self-supervised pre-training and a suitable architecture, often based on a transformer model with encoder and decoder components. This step involves a reconstruction task where the model reconstructs masked parts of the data.
@@
    Fine-tuning: The pre-trained FM is customized for various downstream tasks with minimal labeled data by attaching a task-specific decoder head and performing a few training iterations.
@@
FM development is generally computationally expensive, involving pre-training on datasets of tens of terabytes and models with hundreds of millions or even billions of parameters [47]. However, once developed, FMs can be readily adapted to multiple downstream tasks via fine-tuning with minimal labeled data. Thus, the cost of model pre-training can be spread across numerous applications. These economies of scale make it feasible to employ AI for niche applications, for which developing a dedicated traditional AI model might be prohibitively expensive or technically infeasible.
FM training encodes underlying patterns from the supplied data in model parameters. Successful examples of FMs to date include text-based large language models (LLMs) [47] and weather forecast models [48], in which these patterns are linguistic and physics-driven temporal-spatial relationships, respectively. LLMs have been shown to be applied successfully to certain power system problems [14]. By contrast, our focus here is on a multi-modal GridFM able to handle multiple data modalities intrinsic to power grid operations. Modalities of interest include network topologies, sensor data from phasor measurement units (PMUs), SCADA systems, advanced metering infrastructure (AMI), load, generation, weather, and market data.
@@ 3.2 Salient advantages
We next review salient features of FMs that hold significant potential for enhancing power grid applications.
@@
In transformer models, a token is defined as the smallest fundamental unit of data (e.g., a character sequence or image patch). FMs excel at predicting missing tokens based on context, enabling tasks such as sentence completion, programming assistance, and question answering in the text domain [49]. This capability extends beyond text to fields like time-series, audio, or video, where FMs predict next or missing values [50].
@@
FMs can emulate physics-based simulations with impressive accuracy and efficiency, potentially combined with physics-informed machine learning [28]. The efficiency gains originate from AI models directly mapping inputs to outputs, unlike traditional physics-based simulations that solve partial differential equations iteratively. Recent AI models [48] emulate numerical weather prediction models [52] with comparable forecasting skills but up to 4–5 orders of magnitude lower computational cost [53] during inference. These game-changing advances improve the spatial and temporal resolution of forecasts and, combined with AI techniques such as diffusion-based models [54], facilitate the creation of large ensembles for uncertainty estimation. Another salient advantage of using AI for emulating simulations lies in the fact that no a-priori assumptions are required but rather the FM can learn the full physics directly from the data or observations. We argue that exactly these characteristics of FMs are crucial for bridging the computational gaps shown in Figure 1.
@@ 3.3 Hurdles to implementation
FMs, like other deep learning models, are data hungry, which raises the question of the sufficiency of available data. For example, the Llama3 LLM, with 70 billion parameters, is trained on 15 trillion tokens of text [55]. By way of comparison, the power sector globally generates multiple times that volume daily [56]. Considering the groundbreaking capabilities of large language models in open-ended systems where context and semantics are not consistently described by mathematical equations, there are reasons for optimism about this technology’s potential for the power grid. However, the question remains whether data sets with the right diversity and quality exist or can be assembled so that the full spectrum of scenarios and situations are covered to encode the desired information in model parameters via a suitable training approach.
@@
Data accessibility is another potential hurdle to overcome. Grid data are heterogeneous and distributed, are owned by multiple entities, and have special security and privacy measures. A potential solution to this barrier includes privacy-preserving federated learning [57] and secure enclave technologies [58]. Such approaches may enable multiple data owners to train models collaboratively without sharing their local data with a central server, thus preserving privacy and complying with data protection regulations. Additional schemes like differential privacy and homomorphic encryption can be integrated into federated learning frameworks [59], ensuring that sensitive information remains protected during training. Secure enclaves, in contrast, create a trusted execution environment for secure processing of sensitive and proprietary data during FM training. Moreover, simulated data generated from real and synthetic systems can complement and provide scenarios under various conditions, such as extreme weather conditions, component failures, and cascading failure events [13]. However, it is important to consider that electrical networks evolve for various reasons, such as network reinforcement, which changes topologies and operating conditions, or new business models (e.g., local energy communities), which change the operating profiles of assets connected to the network, requiring continual learning of the FM.
@@
Notwithstanding the scenario in which all necessary data, whether synthetic or real, can be cleared and collected for use, there still remains the age-old question of whether the data and, consequently, a data-driven model can be trusted. In general, power system operators prefer interpretable models over black-box models so that the root causes of choices by advanced tools can be better understood, and heuristic solutions can be developed. While the debate on the trustworthiness of AI will continue, there is evidence that AI models can be as trustworthy as first-principle physics-based models, for specific applications [60]. Referring back to the recent breakthroughs of AI emulators for weather [61], it was demonstrated that AI could indeed encode the actual dynamics of certain physical systems and generalize to unseen events in ways that allow it to outperform a physics-based simulation-based model [62]. Moreover, standardized formal verification methods are required for AI. ISO already developed the first family of standards in that direction, e.g., ISO/IEC 24029-2 [63].
The European Union’s recently published AI Act [64] defines two concepts to boost trust in AI-based systems and data sharing: a) testing and experimentation facilities that combine physical and virtual environments to evaluate and certify their latest AI-based software and hardware, and b) the data spaces concept to remove barriers to data sharing, such as interoperability, trust, and privacy, while keeping sovereignty and full control of data.

### v2

# Foundation Models for the Electric Power Grid

@@
Foundation Models for the Electric Power Grid
Hendrik F. Hamann§* Blazhe Gjorgiev* Thomas Brunschwiler* Leonardo S. A. Martins* Alban Puech* Anna Varbella* Jonas Weiss* Juan Bernabe-Moreno Alexandre Blondin Massé Seong Choi Ian Foster Bri-Mathias Hodge†, Rishabh Jain Kibaek Kim Vincent Mai François Mirallès Martin De Montigny Octavio Ramos-Leaños Hussein Suprême Le Xie El-Nasser S. Youssef Arnaud Zinflou Alexander Belyi Ricardo J. Bessa Bishnu Prasad Bhattarai Johannes Schmude Stanislav Sobolevsky
@@
§ Lead contact: hehamann.hh@gmail.com
* These authors contributed equally to this work as main contributors.
† These authors contributed equally to this work as significant contributors.
‡ These authors contributed equally to this work.
Keywords
Summary
Foundation Models (FMs) currently dominate news headlines. They employ advanced deep learning architectures to extract structural information autonomously from vast data-sets through self-supervision. The resulting rich representations of complex systems and dynamics can be applied to many downstream applications. Therefore, advances in FMs can find uses in electric power grids, challenged by the energy transition and climate change. This paper calls for the development of FMs for electric grids. We highlight their strengths and weaknesses amidst the challenges of a changing grid. It is argued that FMs learning from diverse grid data and topologies, that we call GridFMs, could unlock transformative capabilities, pioneering a new approach in leveraging AI to redefine how we manage complexity and uncertainty in the electric grid. Finally, we discuss a practical implementation pathway and road map of a GridFM–v0, a first GridFM for power flow applications based on graph neural networks, and explore how various downstream use cases will benefit from this model and future GridFMs.
@@ 1 Introduction
Figure 1: The energy transition, aging infrastructure, cybersecurity challenges, and climate change greatly increase complexity and uncertainty in operating, controlling, and planning the power grid. This is creating a widening gap between existing computational capabilities and the evolving needs of the electric power industry.
@@
In this context, it is clear that the operation, control, and planning of power systems will soon be pushed to their limits. As depicted in Figure 1, new computational methods and approaches are needed, capable of better tackling the challenges presented by increased uncertainty and complexity. Machine Learning (ML) and Artificial Intelligence (AI) methods have shown promise for such purposes in a wide spectrum of industries, with significant breakthroughs in computer vision [8], natural language processing [9], and intelligent control [10]. Such approaches rely on statistical algorithms to learn from massive data, often enabling them to execute tasks (e.g., regression, classification, detection, segmentation, forecasting, and generation) without human supervision.
@@
Yet, despite significant progress in the application of AI/ML methods to power systems [11, 12, 13], wider adoption has been impeded, in particular, by a lack of readily available training data and the limited transferability of developed models to adjacent applications and system configurations [14]. This is where, we argue, Foundation Models (FM) have an important role to play. FMs are advanced AI models developed through self-supervised learning, most often based on transformer architectures, that generalize across various tasks after initial training on large data-sets, enabling efficient adaptation to specific applications with minimal labeled data. Text-based FMs like GPT-4 [15], encoding deep relationships in text, have enabled a wide range of new applications. Similarly, we anticipate that FMs for power grids, trained on the diverse data that characterizes the power grid, will enable a broad set of new use cases.
@@
In this article, we advocate for developing FMs for power grids: what we refer to as GridFMs. We explain why we believe in their potential, highlighting their anticipated strengths and weaknesses in light of the challenges posed by the evolving power grid. We introduce a practical road map for implementing GridFM–v0, an initial version of a GridFM for power flow applications, and discuss our vision for future GridFMs, pioneering a new approach in leveraging AI to manage complexity and uncertainty in the power grid.
@@
In the following, we first highlight, in Section 2, the looming power grid challenges associated with the energy transition, which are further aggravated by aging infrastructure, swelling cybersecurity, and climate change concerns. We show how, by increasing the complexity and uncertainty of the power grid, these drivers are creating a significant gap in computational capabilities, which we propose to overcome with FMs. We then review in Section 3 the strengths and weaknesses of FMs and propose in Section 4 the near-term and long-term development phases of GridFMs that we argue will overcome current limitations and enable scenarios, which are currently unattainable.
@@ 2.1 Distributed and renewable energy sources
While a single coal plant can generate hundreds of megawatts predictably and reliably, large numbers of weather-dependent renewable energy resources like wind and solar are required to produce the same power output. Current methods to forecast 1-day or 1-hour ahead still show up to 3% to 10% forecast error, respectively [16], depending on the energy source. Additionally, Distributed Energy Resources (DERs) like battery storage and electric vehicles are operated by independent power management systems and are driven by user patterns with intrinsic uncertainties. Dealing with the added uncertainty from reverse DER power flows, multipoint injections, and weather-induced fluctuations requires new voltage regulation methods, protection schemes, and advanced (optimal) power flow analysis. The numbers of controllable and dispatchable devices, as well as balancing nodes are expected to grow exponentially [17] with additional thousands to millions to be added in the near future. Managing the associated variability and uncertainty will require advanced stochastic optimization, scenario reduction, and sampling-based approaches, which are computationally more intensive than currently used deterministic methods, especially for real-time demand-supply balancing.Furthermore, as the net load of DER-rich distribution branches seen by transmission operators becomes more variable and less predictable, effective grid management and planning will become impossible without effective co-simulation frameworks that integrate transmission and distribution analysis. Yet the computational cost of such simulations, at least with existing technologies, is extremely high.
@@ 2.2 Everything inverter
For the future power grid, inverters will play a crucial role in maintaining low cost at the grid’s edge. However, with increased inverter proliferation, established control schemes and grid stability are challenged, while a more deregulated environment is created. With more inverters connected, grid inertia – which stabilizes phase, voltage, and frequency during transient events and which is mostly tied to large synchronous generators – decreases [18], unless more grid-forming converters are introduced. Because inverters are controlled by software, they can quickly adjust power or disconnect, causing additional issues like sub-synchronous resonances and even trip large generators [19].
[…]
To further account for inherent network uncertainties and those introduced by Inverter-Based and Distributed Energy Resources (IBRs, DERs), stochastic ensembles of the problems must be simulated. Altogether, such an approach is computationally very expensive and, with current approaches, impractical [25]. Thus, the community mostly relies on offline simulators like PS CAD [26] or ParaEMT [23] while starting exploring physics-informed neural networks (PINNs) to reduce EMT simulation time [27].
@@ 2.3 Changes to demand and weather patterns
Long-term models like the Regional Energy Deployment System [34] are used for resource, infrastructure, and strategic planning. Uncertainty and variability at longer time scales are significantly driven by climate policy, technology adoption, and regulatory changes. Instances of this include the combustion engine phase-out by 2035 in Europe – which resulted in a massive increase in the long-term electricity demand forecast – or the US Federal Energy Regulatory Commission’s rules on planning and covering the cost of the power grid [35]. To improve model accuracy, additional information that captures demographic changes and changes in the distribution of wealth, which lead to different societal and individual behaviors (e.g., heating, cooling, electromobility), as well as climate change-induced environmental shifts must be integrated. Therefore, the capacity to combine heterogeneous data sources, including information in textual format, is a fundamental requirement to improve predictability.
@@ 2.4 Aging infrastructure
Moreover, the lowering of grid inertia through IBR integration (see Section 2.2) is further amplified by the retirement of old synchronous generators. The more frequent use of synchronous generators for regulating purpose bring them more often in operating zones where mechanical or thermal stresses are significant, and thus accelerate the aging of the generators. IBR related switching transients also disrupt existing protection relay schemes [40] and increase grid behavioral uncertainty.
@@ 3 Foundation models
Having identified the growing gaps between computational capabilities and industry needs amidst the energy transition, we now introduce FMs, laying the groundwork for our vision of GridFMs.
@@ 3.1 The foundation model approach
To understand the FM approach, it is helpful to review the evolution of AI and ML as depicted in Figure 2. Early hand-crafted symbolic expert systems, from the 1960s onwards, were limited and brittle. The 90s brought big data and general-purpose ML that, however, still often depend on task-specific hand-crafted features. In 2012, deep learning, enabled by increased compute performance, started to disrupt the AI world [45], but was more often than not used to train models on large annotated data-sets in a supervised way, making them task-specific. Recently, FMs have emerged. They learn from data through self-supervision and large content-rich data-sets [46] and have proven to generalize across many applications. FMs draw their potential from being able to train on very large data-sets, which have not been annotated, thus are more easily available than labeled data-sets, as required for supervised learning. The FM approach is a three-step process, as shown in Figure 3:
@@
    Pre-training: A general-purpose FM is developed through self-supervised pre-training and a suitable architecture, often based on a transformer model with encoder and decoder components. This step involves a reconstruction task where the model reconstructs masked parts of the data.
@@
    Fine-tuning: The pre-trained FM is customized for various downstream tasks and applications with minimal labeled data by attaching a task-specific decoder head and performing a few training iterations.
@@
The development of FMs involves extensive pre-training on data-sets of tens of terabytes and models with hundreds of millions or even billions of parameters [47]. However, in contrast to what one might assume, even bespoke AI models can be very complex and expensive to train. For example, the non-foundation and application-specific weather AI models, Graphcast (for forecasting) or CorrDiff (for downscaling) were trained on 32 TPU devices [48] and on 128 A100 GPUs [49], respectively.
The key advantage of FMs is that once developed, they can be readily adapted to multiple downstream applications via fine-tuning with minimal labeled data. Thus, the cost of model pre-training can be spread across numerous applications. These economies of scale make it feasible to employ AI for niche applications, for which developing a dedicated traditional AI model might be prohibitively expensive or technically infeasible.
FM pre-training encodes underlying patterns from the supplied data in model parameters. Successful examples of FMs to date include text-based Large Language Models (LLMs) [47] but also weather forecast models [50], in which these patterns are linguistic and physics-driven temporal-spatial relationships, respectively. LLMs have been shown to be applied successfully to certain power system problems [14]. By contrast, our focus here is on GridFMs that go beyond text, that will eventually be able to handle multiple data modalities intrinsic to power grid operations. Modalities of interest include network topologies, sensor data from Phasor Measurement Units (PMUs), alarms, SCADA systems, Advanced Metering Infrastructure (AMI), load, generation, weather, and market data.
@@ 3.2 Salient advantages
We next review salient features of AI in general and FMs in particular that hold significant potential for enhancing power grid applications.
@@
In transformer models, a token is defined as the smallest fundamental unit of data (e.g., a character sequence or image patch). FMs have been shown to excel at predicting missing tokens based on context, enabling tasks such as sentence completion, programming assistance, and question answering in the text domain [51]. This capability extends beyond text to fields like time-series, audio, or video, where FMs predict next or missing values [52].
@@
AI models in general, including FMs can emulate physics-based simulations with impressive accuracy and efficiency, potentially combined with physics-informed machine learning [27]. The efficiency gains originate from AI models directly mapping inputs to outputs, unlike traditional physics-based simulations that solve partial differential equations iteratively. Recent AI models [54] emulate numerical weather prediction models [55] with comparable forecasting skills but up to 4–5 orders of magnitude lower computational cost [56] during inference. These game-changing advances improve the spatial and temporal resolution of forecasts and, combined with AI techniques such as diffusion-based models [57], facilitate the creation of large ensembles for uncertainty estimation. Another salient advantage of using AI for emulating simulations lies in the fact that no a-priori assumptions are required, but rather the AI model can learn the full physics directly from the data or observations. We argue that exactly these characteristics of AI in general and FMs in particular are crucial for bridging the computational gaps shown in Figure 1.
@@ 3.3 Hurdles to implementation
FMs, like other deep learning models, are data hungry. For example, the Llama3 LLM, with 70 billion parameters, is trained on 15 trillion tokens of text [58]. By way of comparison, the power sector globally generates multiple times that volume daily [59]. Considering the groundbreaking capabilities of large language models in open-ended systems where context and semantics are not consistently described by mathematical equations, we are confident about this technology’s potential for the power grid. Diverse and high-quality data-sets required for training grid FMs will, though, still have to be assembled to cover the full spectrum of scenarios and situations. Addressing transmission grid problems first, where data is available in a more standardized form, e.g., PMU or power flow data will help demonstrate the value of FMs for the grid. This will create incentives for data at the distribution level, which currently is much more heterogeneous and of less quality [60], to be collected, curated, and made available in suitable formats for use with AI and FMs at a larger scale.
@@
Data accessibility is another hurdle to overcome. Grid data are heterogeneous and distributed, are owned by multiple entities, and have special security and privacy measures. A potential solution to this barrier includes privacy-preserving federated learning [61] and secure enclave technologies [62]. Such approaches may enable multiple data owners to train models collaboratively without sharing their local data with a central server, thus preserving privacy and complying with data protection regulations. Additional schemes like differential privacy and homomorphic encryption can be integrated into federated learning frameworks [63], ensuring that sensitive information remains protected during training. Secure enclaves, in contrast, create a trusted execution environment for secure processing of sensitive and proprietary data during FM training. Moreover, simulated data from real and synthetic systems can complement and provide scenarios under various conditions, such as extreme weather conditions, component failures, and cascading failure events [13]. As discussed further in Section 4, GridFMs might be developed by synthetic data, which are then updated and verified with proprietary data behind the user’s firewall. However, it is important to consider that electrical networks evolve for various reasons, such as network reinforcement, which changes topologies and operating conditions, or new business models (e.g., local energy communities), which change the operating profiles of assets connected to the network, requiring continual learning of the FM.
@@
Notwithstanding the scenario in which all necessary data, whether synthetic or real, can be cleared and collected for use, there still remains the age-old question of whether the data and, consequently, a data-driven model can be trusted. In general, power system operators prefer interpretable models over black-box models so that the root causes of choices by advanced tools can be better understood, and heuristic solutions can be developed. While the debate on the trustworthiness of AI will continue, there is evidence that AI models can be as performant as first-principle physics-based models, at least for specific applications [64]. Performance alone however does not seem to be necessarily a sufficient condition for trust because even conventional solvers can have stability and convergence challenges [65], which we have found ways to deal with through decades of usage.
Referring back to the recent breakthroughs of AI emulators for weather, it was demonstrated that AI could indeed encode the actual dynamics of certain physical systems and generalize to unseen events in ways that allow it to have comparable performances to a physics-based, simulation-based model [66]. In some cases, e.g. for an AI power flow emulator, which is discussed in Section 4, the AI results can be immediately verified by using the physics equations. In other cases, where such ad-hoc verification is not possible, a consistent way is required to evaluate the model for out-of-sample scenarios. In the longer-term a set of standards must be developed as to what performances must be achieved for a given out-of sample scenario. Towards that end, ISO has already developed the first family of standards, e.g., ISO/IEC 24029-2 [67]. The European Union’s recently published AI Act [68] defines two concepts to boost trust in AI-based systems and data sharing: a) testing and experimentation facilities that combine physical and virtual environments to evaluate and certify their latest AI-based software and hardware, and b) the data spaces concept to remove barriers to data sharing, such as interoperability, trust, and privacy, while keeping sovereignty and full control of data.